OpenStudy (anonymous):

find the coordinates of the missing endpoint given that P is the midpoint of NQ. N(5,4) P(5,2)

OpenStudy (anonymous):

\[\frac{5+x}{2}=5\]\[ 5+x=5*2\]\[x=10-5\]\[x=5\]\[\frac{4+y}{2}=2\]\[4+y=2*2\]\[y=4-4\]\[y=0\]If you like my answer,click good answer,TY

OpenStudy (anonymous):

midpoint =\[\left( (x _{1}+x _{2})/2,(y _{1}+y _{2})/2) \right)\] given N(5,4) and midpoint (5,2) then \[(5,2)=\left(( 5+x _{2})/2,(2+y _{2})/2) \right)\] then the rest follows from above from Nancy Lam... Q(5,0)